Solve the equation by hand. List answers as exact values, not decimal approximations.

4^{x}\cdot 2^{x^{2}}=16^{2}

The correct answer is: -4,2

Just not sure how to do the problem!
Answer #1

Solucd; 4 t = 162 22(*) - 2x² 2 = 2402) 22x.2012 = 28 - 2x+x² - > 2x+x²=8 22 +22-8 =O x2 +48-2x-8=0 >c(x+4)- 2 x +4)=0 (Oc+4)